Determine the angle which the string would make with the horizontal,whan weights of the blocks A and B are such that


(a)WA=WB=W and


(b)WA=W,WB=2W.

1
Expert's answer
2022-01-26T12:03:02-0500


1. Analyzing vertical forces where string 1 and string 2 meet, we find


"T_1 =\frac{W_A}{cos\phi}" N


2. Looking at the horizontal forces at that point leads to 


"T_2 =T_1Sin \theta" N


3.The angle of the string 


"\theta =tan^{-1}\frac{T_2}{T_1}"